OpenStudy (anonymous):

Combinatorics: In how many variations can you create a team of 7 people from the following groups: A: 6 People B: 5 People C: 4 People D: 3 People E: 2 People When the team has at least 3 people from group B.

OpenStudy (anonymous):

Would this be correct?:\[\left(\begin{matrix}5 \\ 3\end{matrix}\right)\times \left(\begin{matrix}17 \\ 4\end{matrix}\right)=23800\]

OpenStudy (dangerousjesse):

Yes, that looks right to me.

OpenStudy (anonymous):

Right, and what about when 4 are taken from B, and 5 people from B?

OpenStudy (anonymous):

Oh hold on, I though it said \(\dbinom{10}4\)... Never mind!

OpenStudy (anonymous):

I think they're taken into account in the "17 above 4", since they asked for at least 3. I'm not 100% sure you have to include these as well.

OpenStudy (anonymous):

No you're right, I misread your computation.

OpenStudy (anonymous):

Thank you both for the confirmations.